OpenStudy (natalie1234):

I'll Give A Medal! If f(x) varies directly with x, and f(x) = 12 when x = 8, write the direct linear variation equation

OpenStudy (the_fizicx99):

f(x) = kx OR y = kx (f(x) is another way of saying y) 12 = 8k /8 /8 1.5 12 = 8*1.5 12 = 12

OpenStudy (natalie1234):

thanks! :)